OpenStudy (anonymous):

30 mi / h = ____ ft / min

OpenStudy (anonymous):

can anyone help?

OpenStudy (aravindg):

Use the fact that: 1 mi/h =88 feet / minute

OpenStudy (anonymous):

400 feet in 52.4 seconds im also stuck on this

OpenStudy (aravindg):

^Sorry didnt get you.

OpenStudy (anonymous):

express as a unit rate

OpenStudy (anonymous):

400 feet in 52.4 seconds

OpenStudy (aravindg):

400/52.4=x/1 If thats what you meant. Solve for x.

OpenStudy (anonymous):

Got a 80% on my test thank you so much i need a personal tutor want to help me out ?

OpenStudy (aravindg):

I can be a mentor.

OpenStudy (anonymous):

8134108082

OpenStudy (anonymous):

hit me up

OpenStudy (aravindg):

Sorry I am in a different country.

OpenStudy (aravindg):

You could post your questions here.

OpenStudy (anonymous):

okay cool stay on i have some stuff in need of

OpenStudy (anonymous):

A 3” x 5” photo is enlarged so that the length of the new photo is 7 inches. Find the width of the new photo

OpenStudy (aravindg):

Posrt your questions separately not in the same question.

OpenStudy (kenljw):

Creation of a conversion factor 1 mi = 4,280 ft 1 hr = 60 min 1 mi/hr = 4,280/60 ft/min = 71 1/3 ft/min 30 mi / hr = 30 (71 1/3) ft/min = 2140 ft/min